OpenStudy (anonymous):

write algebraic expression for sec(arcsin(x/sqrtx^2+4))

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sec\left(\arcsin\frac{x}{\sqrt{x^2+4}}\right) }\) like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sec\left(A\right)= \frac{1}{\cos(A)}=\frac{1}{\sqrt{1-\sin^2(A)}}}\)

OpenStudy (solomonzelman):

And we can use that here

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sec\left(\arcsin\frac{x}{\sqrt{x^2+4}}\right)=\frac{1}{\sqrt{1-\sin^2\left(\arcsin\dfrac{x}{\sqrt{x^2+4}}\right)}} }\)

OpenStudy (solomonzelman):

The thing is, that: \(\color{red}{\sin\left(\arcsin B\right)=B}\)

OpenStudy (solomonzelman):

And so, \(\color{red}{\sin~^m~\left(\arcsin B\right)=B~^m}\)

OpenStudy (anonymous):

sorry computer is super slow right now

OpenStudy (solomonzelman):

All that matters is whether you are able to follow what I am trying to get you to do or not.... so are you with me?

OpenStudy (anonymous):

ummm in a way. but the directions say that i must use the right triangle method.

OpenStudy (solomonzelman):

Ok, you didn't mention that... sadly. But, still can you give me the answer based on what I have suggested, and then we will do this your way

OpenStudy (anonymous):

My bad. i thought there was only one method to solve this.

OpenStudy (solomonzelman):

(I worked this out the answer is pretty simple)

OpenStudy (anonymous):

all right so then it is\[\frac{ 1 }{\sqrt{1-\frac{ x }{\sqrt{x^2+4} }} }\]

OpenStudy (solomonzelman):

no not quite

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle \sin^{\rm P}\left(\arcsin {\bf Q}\right)={\bf Q}^{\rm P} }\)

OpenStudy (solomonzelman):

and that follows logically, just knowing the rules of exponents and knowing the fact that \(\Large\color{black}{ \displaystyle \sin\left(\arcsin {\bf Q}\right)={\bf Q} }\)

OpenStudy (solomonzelman):

So, \(\large\color{black}{ \displaystyle \frac{1}{\sqrt{1-\sin^{\color{red}{\LARGE 2}}\left(\arcsin\dfrac{x}{\sqrt{x^2+4}}\right)}}=?}\)

OpenStudy (anonymous):

umm \[\frac{ 1 }{ \sqrt{1-\frac{ x^2 }{ (x+4)^2 }} }\]

OpenStudy (anonymous):

right?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1}{\sqrt{1-\sin^2\left(\arcsin\dfrac{A}{B}\right)}}=\frac{1}{\sqrt{1-\frac{A^2}{B^2}}} }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1}{\sqrt{1-\sin^2\left(\arcsin\dfrac{x}{\sqrt{x^2+4}}\right)}}=\frac{1}{\sqrt{1-\dfrac{(x)^2}{(\sqrt{x^2+4})^2}}}}\)

OpenStudy (anonymous):

is that the final

OpenStudy (solomonzelman):

so you will get \(\large\color{black}{ \displaystyle \frac{1}{\sqrt{1-\dfrac{x^2}{|x^2+4|}}}}\) absolute value is purely technical, but we assume the x is real so it is anyways positive. (and no you can simplify)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1}{\sqrt{1-\dfrac{x^2}{x^2+4}}}}\)

OpenStudy (anonymous):

oh ok.

OpenStudy (solomonzelman):

Hint to simplify: \(\large\color{black}{ \displaystyle \frac{1}{\sqrt{\dfrac{x^2+4}{x^2+4}-\dfrac{x^2}{x^2+4}}}}\)

OpenStudy (anonymous):

Can we do the right triangle method?

OpenStudy (anonymous):



OpenStudy (anonymous):

this is the right triangle method. i have to solve for A. so i will use the pythagoras theory

OpenStudy (anonymous):

So.... A^2+x^2=(sqrtx^2+4)^2

OpenStudy (freckles):

yep and solve for A

OpenStudy (freckles):

 then you find sec(u)

OpenStudy (anonymous):

yeah but i dont know how to set it up

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

you already did just solve for A

OpenStudy (freckles):

\[A^2+x^2=(\sqrt{x^2+4})^2 \\ A^2+x^2=x^2+4 \\ \\ \text{ subtract } x^2 \text{ on both sides } \\ A^2=4\]